### v2

@@ 4 Convergence Analysis of The P-BCD Algorithm
(23) $\lim_{k\to\infty}\left\|[H_{i}^{k}]^{\top}[{X}_{i}^{k}]-{[{X}_{i}^{k}]}^{\top}[H_{i}^{k}]\right\|_{F}=0.$
Since $\mathcal{Z}^{\tilde{k}}\to\bar{\mathcal{Z}}$ as $\tilde{k}\in\mathcal{K}$ approaches $\infty$, we have $[{P}_{i}^{\tilde{k}}]\to[\bar{P}_{i}]$, $[Q_{1}^{\tilde{k}}]\to[\bar{Q}_{1}]$, as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. Also, we have
$\|[Q_{2}^{\tilde{k}}]-[\bar{Q}_{2}]\|_{F}$ $\leq\|[Q_{2}^{\tilde{k}}]-[\tilde{Q}_{2}^{\tilde{k}}]\|_{F}+\|[\tilde{Q}_{2}^{\tilde{k}}]-[\bar{Q}_{2}]\|_{F}$
$\leq c_{1}\|[X_{1}^{\tilde{k}+1}]-[X_{1}^{\tilde{k}}]\|_{F}+\|[\tilde{Q}_{2}^{\tilde{k}}]-[\bar{Q}_{2}]\|_{F}$
@@
$\|[Q_{3}^{\tilde{k}}]-[\bar{Q}_{3}]\|_{F}$ $\leq\|[Q_{3}^{\tilde{k}}]-[\tilde{Q}_{3}^{\tilde{k}}]\|_{F}+\|[\tilde{Q}_{3}^{\tilde{k}}]-[\bar{Q}_{3}]\|_{F}$
$\leq c_{1}\|[X_{1}^{\tilde{k}+1}]-[X_{1}^{\tilde{k}}]\|_{F}+c_{1}\|[X_{2}^{\tilde{k}+1}]-[X_{2}^{\tilde{k}}]\|_{F}+\|[\tilde{Q}_{3}^{\tilde{k}}]-[\bar{Q}_{3}]\|_{F},$
where $[\tilde{Q}_{2}^{\tilde{k}}]=\big{(}[\mathcal{G}^{\tilde{k}}]\times_{1}[X_{1}^{\tilde{k}}]\times_{3}[X^{\tilde{k}}_{3}]\big{)}_{[(2)]}$ and $[\tilde{Q}_{3}^{\tilde{k}}]=\left([\mathcal{G}^{\tilde{k}}]\times_{1}[X^{\tilde{k}}_{1}]\times_{2}[X^{\tilde{k}}_{2}]\right)_{[(3)]}$. Since $[\tilde{Q}_{i}^{\tilde{k}}]\to[\bar{Q}_{i}]$, $i=2,3$, by Theorem 4.1 (iii), we have $[Q_{i}^{\tilde{k}}]\to[\bar{Q}_{i}]$, $i=2,3$, as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. Hence, $[H_{i}^{\tilde{k}}]\to[\bar{H}_{i}]$, $i=1,2,3$, as ${k}\in\mathcal{K}$ approaches $\infty$. By (22) and (23), we have (15b) and (15c) hold. Since Stiefel manifold is a compact set, we also have (15d) holds.
@@ 5 Application to HSI Denoising and Destriping
To denoise the HSI via the low-rank tensor regularization, we extract global, local, and nonlocal tensors that may have low-rank features via the block extraction operator $\mathbf{R}$ and impose the low-rankness on the tensors. In the following, we first introduce the block extraction for global, local, and nonlocal tensors, i.e., the choices of $\mathbf{R}$, and the sparsity enhancement on low-rank regularization for removing Gaussian noise, i.e., the choice of $\psi$. Then we introduce the tensor $\ell_{2,p}$ norm for removing sparse noises with linear structures, i.e., the choice of $\varphi$. Lastly, we propose the MLTL2p method using a multi-scale low-rank tensor regularization with a two-phase scheme to increase the effectiveness and efficiency of model (1) and Algorithm 1.
@@ 5.1 Block extraction
According to the spectral correlation and spatial nonlocal self-similarity of HSIs, a clean HSI can be approximated using global, local, and nonlocal low-rank tensors [26, 28]. These tensors differ in their extraction operators $\mathbf{R}$, which we later denote as $\mathbf{R}_{g}$ for global, $\mathbf{R}_{l}$ for local, and $\mathbf{R}_{nl}$ for nonlocal.
We begin by defining several types of block extraction operators and their transpose operators. Given an HSI $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we can divide it into a total number of $N$ overlapping or non-overlapping blocks of size $r_{1}\times r_{2}\times r_{3}$, $r_{i}\leq I_{i}$. We define the operator that extracts the $l$-th block as $\mathbf{B}_{l}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$
$\mathbf{B}_{l}(\mathcal{L}):=\mathcal{L}\times_{1}(U_{1}^{(l)})^{\top}\times_{2}(U_{2}^{(l)})^{\top}\times_{3}(U_{3}^{(l)})^{\top},$
where $U_{i}^{(l)}\in\mathbb{R}^{I_{i}\times r_{i}}$ is a binary matrix such that $\mathbf{B}_{l}(\mathcal{L})$ is exactly the $l$-th local block of $\mathcal{L}$, and $(U_{i}^{(l)})^{\top}U_{i}^{(l)}=\mathbf{I}_{r_{i}}$. The transpose of $\mathbf{B}_{l}$ can be defined by $\mathbf{B}_{l}^{\top}:\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$
$\mathbf{B}_{l}^{\top}(\mathcal{Y}):=\mathcal{Y}\times_{1}U_{1}^{(l)}\times_{2}U_{2}^{(l)}\times_{3}U_{3}^{(l)},$
which satisfies $\langle\mathcal{Y},\mathbf{B}_{l}(\mathcal{L})\rangle_{F}=\langle\mathbf{B}^{\top}_{l}(\mathcal{Y}),\mathcal{L}\rangle_{F}$ for any $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\mathcal{Y}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$. For those $N$ independent blocks, we can stack them together and form a new operator $\mathbf{B}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{r_{1}\times r_{2}\times r_{3}\times N}$ such that $[\mathbf{B}(\mathcal{L})]^{(l)}=\mathbf{B}_{l}(\mathcal{L})$.
Depending on the size of the blocks, there are several types of block extraction as follows using the formulation of $\mathbf{B}$:
  • Global extraction $\mathbf{R}_{g}$: $\mathbf{B}$ with $r_{i}=I_{i}$, $i=1,2,3$;
  • Local extraction $\mathbf{R}_{l}$: $\mathbf{B}$ with $r_{i}<I_{i}$, $i=1,2,3$;
  • Full band extraction $\mathbf{R}_{f}$: $\mathbf{B}$ with $r_{i}<I_{i}$, $i=1,2$ and $r_{3}=I_{3}$.
@@
Next, we formulate another type of block extraction operators based on spatial nonlocal self-similarity of HSIs. After dividing the HSI into $N$ blocks using $\mathbf{B}_{l}$, we apply block matching to find similar blocks and then stack them into a fourth order nonlocal similar tensor.
Taking FBBs as an example, for the $j$-th FBB, we search within a local window for a total of $m_{2}$ FBBs that are mostly similar to the reference FBB based on Euclidean distance. Then the $j$-th nonlocal similar sub-tensor of order 3 of $\mathcal{L}$, denoted as $\mathbf{R}_{j}(\mathcal{L})$, can be formed by unfolding all the nonlocal similar blocks in the $j$-th group and then stacking them together. We define the Casorati matrix (a matrix whose columns are vectorized bands of the HSI) of the $l$-th FBB as follows
$\tilde{B}_{l}:=\operatorname{reshape}(\mathbf{B}_{l}(\mathcal{L}),m_{1},m_{3}),$
where $m_{1}=r_{1}r_{2}$ and $m_{3}=I_{3}$. Then the extraction operator of the $j$-th nonlocal similar sub-tensor $\mathbf{R}_{j}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ can be defined by
$\mathbf{R}_{j}(\mathcal{L}):=\operatorname{reshape}([\tilde{{B}}_{l_{j1}}^{\top},\tilde{{B}}_{l_{j2}}^{\top},\dots,\tilde{{B}}_{l_{jm_{2}}}^{\top}]^{\top},m_{1},m_{2},m_{3}),$
where the indices $l_{j1},l_{j1},\dots,l_{jm_{2}}$ refer to the indices of FBBs that belongs to the $j$-th nonlocal similar group and $j=1,2,\dots,N$. An illustrative figure on nonlocal low-rank tensor extraction is shown in Fig. 4. Then the transpose $\mathbf{R}^{\top}_{j}:\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is defined by
(24) $\mathbf{R}^{\top}_{j}(\mathcal{Y}):=\sum_{i=1}^{m_{2}}\mathbf{B}_{l_{ji}}^{\top}\left(\operatorname{reshape}(Y_{:i:},r_{1},r_{2},I_{3})\right),$
which satisfies $\langle\mathcal{Y},\mathbf{R}_{j}(\mathcal{L})\rangle_{F}=\sum_{i=1}^{m_{2}}\langle{Y}_{:i:},\mathbf{B}_{l_{ji}}(\mathcal{L})\rangle_{F}=\langle\mathbf{R}^{\top}_{j}(\mathcal{Y}),\mathcal{L}\rangle_{F}$. Stacking $\mathbf{R}_{j}(\mathcal{L})$ into a fourth order nonlocal similar group tensor, the extraction operator $\mathbf{R}_{nl}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$ is defined as a linear map such that $[\mathbf{R}_{nl}(\mathcal{L})]^{(j)}=\mathbf{R}_{j}(\mathcal{L})$.
@@
Note that the $\mathbf{R}$ formulated above all satisfy the following equality
$\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\mathbf{R}^{\top}\mathbf{R},$
where $\odot$ represents the pointwise multiplication, and each entry of $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents the number of blocks to which the corresponding pixel belongs. For $\mathbf{R}_{g}$, $\mathbf{R}_{l}$ and $\mathbf{R}_{f}$, the weight tensor can be defined by
$\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\sum_{l=1}^{N}\mathbf{B}_{l}^{\top}\mathbf{B}_{l},$
and for $\mathbf{R}_{nl}$, the weight tensor can be defined by
$\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\sum_{j=1}^{N}\mathbf{R}_{j}^{\top}\mathbf{R}_{j}.$
Since we assume that each pixel belongs to at least one low-rank group, we have $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}_{++}$.

### v3

@@ 4 Convergence Analysis of The P-BCD Algorithm
(24) $\lim_{k\to\infty}\left\|[\boldsymbol{H}_{i}^{k}]^{\top}[\boldsymbol{X}_{i}^{k}]-{[\boldsymbol{X}_{i}^{k}]}^{\top}[\boldsymbol{H}_{i}^{k}]\right\|_{F}=0.$
Since $\mathcal{Z}^{\tilde{k}}\to\bar{\mathcal{Z}}$ as $\tilde{k}\in\mathcal{K}$ approaches $\infty$, we have $[\boldsymbol{P}_{i}^{\tilde{k}}]\to[\bar{\boldsymbol{P}}_{i}]$, $[\boldsymbol{Q}_{1}^{\tilde{k}}]\to[\bar{\boldsymbol{Q}}_{1}]$, as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. Also, we have
$\|[\boldsymbol{Q}_{2}^{\tilde{k}}]-[\bar{\boldsymbol{Q}}_{2}]\|_{F}$ $\leq\|[\boldsymbol{Q}_{2}^{\tilde{k}}]-[\tilde{\boldsymbol{Q}}_{2}^{\tilde{k}}]\|_{F}+\|[\tilde{\boldsymbol{Q}}_{2}^{\tilde{k}}]-[\bar{\boldsymbol{Q}}_{2}]\|_{F}$
$\leq c_{1}\|[\boldsymbol{X}_{1}^{\tilde{k}+1}]-[\boldsymbol{X}_{1}^{\tilde{k}}]\|_{F}+\|[\tilde{\boldsymbol{Q}}_{2}^{\tilde{k}}]-[\bar{\boldsymbol{Q}}_{2}]\|_{F}$
@@
$\|[\boldsymbol{Q}_{3}^{\tilde{k}}]-[\bar{\boldsymbol{Q}}_{3}]\|_{F}$ $\leq\|[\boldsymbol{Q}_{3}^{\tilde{k}}]-[\tilde{\boldsymbol{Q}}_{3}^{\tilde{k}}]\|_{F}+\|[\tilde{\boldsymbol{Q}}_{3}^{\tilde{k}}]-[\bar{\boldsymbol{Q}}_{3}]\|_{F}$
$\leq c_{1}\|[\boldsymbol{X}_{1}^{\tilde{k}+1}]-[\boldsymbol{X}_{1}^{\tilde{k}}]\|_{F}+c_{1}\|[\boldsymbol{X}_{2}^{\tilde{k}+1}]-[\boldsymbol{X}_{2}^{\tilde{k}}]\|_{F}+\|[\tilde{\boldsymbol{Q}}_{3}^{\tilde{k}}]-[\bar{\boldsymbol{Q}}_{3}]\|_{F},$
where $[\tilde{\boldsymbol{Q}}_{2}^{\tilde{k}}]=\big{(}[\mathcal{G}^{\tilde{k}}]\times_{1}[\boldsymbol{X}_{1}^{\tilde{k}}]\times_{3}[\boldsymbol{X}^{\tilde{k}}_{3}]\big{)}_{[(2)]}$ and $[\tilde{\boldsymbol{Q}}_{3}^{\tilde{k}}]=\left([\mathcal{G}^{\tilde{k}}]\times_{1}[\boldsymbol{X}^{\tilde{k}}_{1}]\times_{2}[\boldsymbol{X}^{\tilde{k}}_{2}]\right)_{[(3)]}$. Since $[\tilde{\boldsymbol{Q}}_{i}^{\tilde{k}}]\to[\bar{\boldsymbol{Q}}_{i}]$, $i=2,3$, by Theorem 4.1 (iii), we have $[\boldsymbol{Q}_{i}^{\tilde{k}}]\to[\bar{\boldsymbol{Q}}_{i}]$, $i=2,3$, as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. Hence, $[\boldsymbol{H}_{i}^{\tilde{k}}]\to[\bar{\boldsymbol{H}}_{i}]$, $i=1,2,3$, as ${k}\in\mathcal{K}$ approaches $\infty$. By (23) and (24), we have (16b) and (16c) hold. Since Stiefel manifold is a compact set, we also have (16d) holds.
@@ 5 Application to HSI Denoising and Destriping
To denoise the HSI via the low-rank tensor regularization, we extract global, local, and nonlocal tensors that may have low-rank features via the block extraction operator $\mathscr{R}$ and impose the low-rankness on the tensors. In the following, we first introduce the block extraction for global, local, and nonlocal tensors, i.e., the choices of $\mathscr{R}$, and the sparsity enhancement on low-rank regularization for removing Gaussian white noise, i.e., the choice of $\psi$. Then we introduce the tensor $\ell_{2,p}$ norm for removing sparse noise with linear structures, i.e., the choice of $\varphi$. Lastly, we propose the MLTL2p method using a multi-scale low-rank tensor regularization with a two-phase scheme to increase the effectiveness and efficiency of model (1) and Algorithm 1.
@@ 5.1 Block extraction
According to the spectral correlation and spatial nonlocal self-similarity of HSIs, a clean HSI can be approximated using global, local, and nonlocal low-rank tensors [27, 29]. These tensors differ in their extraction operators $\mathscr{R}$, which we later denote as $\mathscr{R}_{g}$ for global, $\mathscr{R}_{l}$ for local, and $\mathscr{R}_{nl}$ for nonlocal.
We begin by defining several types of block extraction operators and their transpose operators. Given an HSI $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we can divide it into a total number of $N$ overlapping or non-overlapping blocks of size $r_{1}\times r_{2}\times r_{3}$, $r_{i}\leq I_{i}$. We define the operator that extracts the $l$-th block as $\mathscr{B}_{l}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$
$\mathscr{B}_{l}(\mathcal{L}):=\mathcal{L}\times_{1}(\boldsymbol{U}_{1}^{(l)})^{\top}\times_{2}(\boldsymbol{U}_{2}^{(l)})^{\top}\times_{3}(\boldsymbol{U}_{3}^{(l)})^{\top},$
where $\boldsymbol{U}_{i}^{(l)}\in\mathbb{R}^{I_{i}\times r_{i}}$ is a binary matrix such that $\mathscr{B}_{l}(\mathcal{L})$ is exactly the $l$-th local block of $\mathcal{L}$, and $(\boldsymbol{U}_{i}^{(l)})^{\top}\boldsymbol{U}_{i}^{(l)}=\mathbf{I}_{r_{i}}$. The transpose of $\mathscr{B}_{l}$ can be defined by $\mathscr{B}_{l}^{\top}:\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$
$\mathscr{B}_{l}^{\top}(\mathcal{Y}):=\mathcal{Y}\times_{1}\boldsymbol{U}_{1}^{(l)}\times_{2}\boldsymbol{U}_{2}^{(l)}\times_{3}\boldsymbol{U}_{3}^{(l)},$
which satisfies
$\langle\mathcal{Y},\mathscr{B}_{l}(\mathcal{L})\rangle_{F}=\langle\mathscr{B}^{\top}_{l}(\mathcal{Y}),\mathcal{L}\rangle_{F},$
for any $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\mathcal{Y}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$. For those $N$ independent blocks, we can stack them together and form a new operator $\mathscr{B}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{r_{1}\times r_{2}\times r_{3}\times N}$ such that $[\mathscr{B}(\mathcal{L})]^{(l)}=\mathscr{B}_{l}(\mathcal{L})$.
Depending on the size of the blocks, there are several types of block extraction as follows using the formulation of $\mathscr{B}$:
  • Global extraction $\mathscr{R}_{g}$: $\mathscr{B}$ with $r_{i}=I_{i}$, $i=1,2,3$;
  • Local extraction $\mathscr{R}_{l}$: $\mathscr{B}$ with $r_{i}<I_{i}$, $i=1,2,3$;
  • Full band extraction $\mathscr{R}_{f}$: $\mathscr{B}$ with $r_{i}<I_{i}$, $i=1,2$ and $r_{3}=I_{3}$.
@@
Next, we formulate another type of block extraction operators, denoted as $\mathscr{R}_{nl}$, based on spatial nonlocal self-similarity of HSIs. After dividing the HSI into $N$ blocks using $\mathscr{B}_{l}$, we apply block matching to find similar blocks and then stack them into a fourth order nonlocal similar tensor.
Taking FBBs as an example, for the $j$-th FBB, we search within a local window for a total of $m_{2}$ FBBs that are mostly similar to the reference FBB based on Euclidean distance. Then the $j$-th nonlocal similar sub-tensor of order 3 of $\mathcal{L}$, denoted as $\mathscr{R}_{j}(\mathcal{L})$, can be formed by unfolding all the nonlocal similar blocks in the $j$-th group and then stacking them together. We define the Casorati matrix (a matrix whose columns are vectorized bands of the HSI) of the $l$-th FBB as follows
$\tilde{\boldsymbol{B}}_{l}:=\operatorname{reshape}(\mathscr{B}_{l}(\mathcal{L}),m_{1},m_{3}),$
where $m_{1}=r_{1}r_{2}$ and $m_{3}=I_{3}$. Then the extraction operator of the $j$-th nonlocal similar sub-tensor $\mathscr{R}_{j}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ can be defined by
$\mathscr{R}_{j}(\mathcal{L}):=\operatorname{reshape}([\tilde{\boldsymbol{B}}_{l_{j1}}^{\top},\tilde{\boldsymbol{B}}_{l_{j2}}^{\top},\dots,\tilde{\boldsymbol{B}}_{l_{jm_{2}}}^{\top}]^{\top},m_{1},m_{2},m_{3}),$
where the indices $l_{j1},l_{j1},\dots,l_{jm_{2}}$ refer to the indices of FBBs that belongs to the $j$-th nonlocal similar group and $j=1,2,\dots,N$. An illustrative figure on nonlocal low-rank tensor extraction is shown in Fig. 4. Then the transpose $\mathscr{R}^{\top}_{j}:\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is defined by
(25) $\mathscr{R}^{\top}_{j}(\mathcal{Y}):=\sum_{i=1}^{m_{2}}\mathscr{B}_{l_{ji}}^{\top}\left(\operatorname{reshape}(\boldsymbol{Y}_{:i:},r_{1},r_{2},I_{3})\right),$
which satisfies $\langle\mathcal{Y},\mathscr{R}_{j}(\mathcal{L})\rangle_{F}=\sum_{i=1}^{m_{2}}\langle{Y}_{:i:},\mathscr{B}_{l_{ji}}(\mathcal{L})\rangle_{F}=\langle\mathscr{R}^{\top}_{j}(\mathcal{Y}),\mathcal{L}\rangle_{F}$. Stacking $\mathscr{R}_{j}(\mathcal{L})$ into a fourth order nonlocal similar group tensor, the extraction operator $\mathscr{R}_{nl}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$ is defined as a linear map such that $[\mathscr{R}_{nl}(\mathcal{L})]^{(j)}=\mathscr{R}_{j}(\mathcal{L})$.
@@
Note that the $\mathscr{R}$ formulated above all satisfy the following equality
$\mathcal{W}_{\mathscr{R}}\odot\operatorname{Id}=\mathscr{R}^{\top}\mathscr{R},$
where $\odot$ represents the pointwise multiplication, and each entry of $\mathcal{W}_{\mathscr{R}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents the number of blocks to which the corresponding pixel belongs. For $\mathscr{R}_{g}$, $\mathscr{R}_{l}$ and $\mathscr{R}_{f}$, the weight tensor can be defined by
$\mathcal{W}_{\mathscr{R}}\odot\operatorname{Id}=\sum_{l=1}^{N}\mathscr{B}_{l}^{\top}\mathscr{B}_{l},$
and for $\mathscr{R}_{nl}$, the weight tensor can be defined by
$\mathcal{W}_{\mathscr{R}}\odot\operatorname{Id}=\sum_{j=1}^{N}\mathscr{R}_{j}^{\top}\mathscr{R}_{j}.$
Since we assume that each pixel belongs to at least one low-rank group, we have $\mathcal{W}_{\mathscr{R}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}_{++}$.